A cannon shoots a projectile horizontally (x) at 50 m/s from the top of a 70 m tall cliff

a. How much time does it take the projectile to reach the ground?

b. What is the range (dx) of the projectile?


1
Expert's answer
2021-05-26T10:35:40-0400

We have given,

"u_x = 50m\/s"

"d= 70m"

a.) "y = u_yt+ \dfrac{1}{2}gt^2"

"70 = \dfrac{1}{2}(9.8)(t)^2"


"t^2 = \dfrac{70 \times 2}{9.8}"


"t = 3.77s"


b.)"Range = u_x \times t"


"= 50 \times 3.77"

"= 188.98m"